OpenStudy (anonymous):

help please....The curve y = ax2 + bx + c is tangent to the line y = x in the origin. It is also tangent to the line y = 2x +3. Determine the function

OpenStudy (anonymous):

f=ax^2 + bx + c The derivative of the curve f'(x)=2ax+b y=x is tangent to the curve in the origin, so at (0,0) The means f(0)=0=a0^2+b0+c=c so c=0 b=1 because at the origin f'(0)=2*0+b=b (that is the slope of the tangent) y=2x+3 is another tangent but we don't know where. this line has slope 2 so we need to find where 2=f'(x) 2=2a+b (b=1) 2=2a+1, solving gives a=1/2 So the curve is f(x)=1/2 x^2+x

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

but where didi the x go in 2=2ax+b